Refer to the accompanying table below According to the Cost-Benefit Principle, how many units of this activity should be carried out? Units of ActivityTotal Cost Total Benefit se $12 $22 $30 536 arded se $6 $12 $20 ok $42 $56 42 $43 6
Answer #1

Activity (Q) should be carried out until Marginal cost (MC) = Marginal benefit (MB) and MB > MC.

MC = Change in total cost (TC) / Change in Q

MB = Change in total benefit (TB) / Change in Q

When Q = 2, TC = $6 and TB = $22.

When Q = 3, TC = $12 and TB = $30.

When Q = 4, TC = $20 and TB = $36.

Therefore,

When Q = 3,

MC = $(12 - 6) / (3 - 2) = $6

MB = $(30 - 22) / (3 - 2) = $8

Therefore MB > MC.

When Q = 4,

MC = $(20 - 12) / (4 - 3) = $8

MB = $(36 - 30) / (4 - 3) = $6

Therefore MB < MC.

Therefore, Q = 3 is optimal activity level.